# Are supernovae driving turbulence in the solar neighborhood?

### 13 Mar 2026 v1

@@ Appendix F Dependence on temperature thresholds
Figure 8: Same as Figure 1, but selecting gas within different temperature ranges: (a) $0<T<10000\,\mathrm{K}$; (b) $200<T<10000\,\mathrm{K}$; (c) $100<T<50000\,\mathrm{K}$.
In this appendix, we investigate how the inferred velocity dispersions depend on the adopted temperature thresholds used to select the H i component in the simulation.
Observationally, neutral atomic hydrogen spans a broad temperature range. The cold neutral medium (CNM) typically occupies $T\sim 25$$250\,\mathrm{K}$, the unstable neutral medium (UNM) $T\sim 250$$4000\,\mathrm{K}$, and the warm neutral medium (WNM) $T\sim 6000$$8000\,\mathrm{K}$ (McClure-Griffiths et al. 2023). In the FRIGG SN runs, Colman et al. (2025) classify the ISM gas phases using temperature thresholds, with the atomic phases occupying temperatures below $\sim 10000\,\mathrm{K}$. In the TIGRESS simulation (Kim and Ostriker 2017), gas phases are defined using temperature thresholds, with neutral gas below $20000\,\mathrm{K}$ and ionized gas above this value. Although molecular hydrogen is present at the lowest temperatures, near those of the CNM, and ionized gas appears at temperatures approaching and exceeding those of the WNM, the total mass over intermediate temperatures is dominated by neutral atomic hydrogen. In their Solar Cirle simulation, Tress et al. (2025) shows that neutral atomic gas dominates the mass distribution particularly between $\sim 100$ and $10000\,\mathrm{K}$ where it exceeds both molecular (H2) and ionized (H+) gas. At lower temperatures ($\lesssim 100\,\mathrm{K}$), the mass is dominated by molecular gas, while at higher temperatures ($\gtrsim 20000,\mathrm{K}$), the ionized gas becomes the dominant component.
@@
Figure 8 shows the distributions of LOS velocity dispersions obtained for different temperature thresholds, for the simulations SN-1.5, TIGRESS-MHD4pc, TURB-7, and TURB-10. Varying the adopted temperature range produces only minor changes in the resulting dispersions. The relative order of the simulations remains unchanged, and the mean values shift by a few tenths of km  s-1 at most. In particular, SN-driven models remain systematically below the observational reference, even when molecular gas is excluded or a fraction of the ionized gas is included, while the large-scale forcing runs continue to exceed it. Our conclusions therefore do not depend sensitively on the adopted temperature range used to approximate the neutral atomic component.

### 31 Mar 2026 v2

@@ Appendix F Dependence on temperature thresholds
Figure 8: Same as Fig. 1 but selecting gas within different temperature ranges: (a) $0<T<10000\,\mathrm{K}$; (b) $200<T<10000\,\mathrm{K}$; (c) $100<T<50000\,\mathrm{K}$.
In this appendix we investigate how the inferred velocity dispersions depend on the adopted temperature thresholds used to select the H i component in the simulation.Observationally, neutral atomic hydrogen spans a broad temperature range. The cold neutral medium (CNM) typically occupies $T\sim 25$$250\,\mathrm{K}$, the unstable neutral medium (UNM) $T\sim 250$$4000\,\mathrm{K}$, and the warm neutral medium (WNM) $T\sim 6000$$8000\,\mathrm{K}$ (McClure-Griffiths et al. 2023). In the FRIGG SN runs, Colman et al. (2025) classify the ISM gas phases using temperature thresholds, with the atomic phases occupying temperatures below $\sim 10000\,\mathrm{K}$. In the TIGRESS simulation (Kim and Ostriker 2017), gas phases are defined using temperature thresholds, with neutral gas below $20000\,\mathrm{K}$ and ionized gas above this value. Although molecular hydrogen is present at the lowest temperatures, near those of the CNM, and ionized gas appears at temperatures approaching and exceeding those of the WNM, the total mass over intermediate temperatures is dominated by neutral atomic hydrogen. In their Solar Cirle simulation, Tress et al. (2025) shows that neutral atomic gas dominates the mass distribution particularly between $\sim 100$ and $10000\,\mathrm{K}$ where it exceeds both molecular (H2) and ionized (H+) gas. At lower temperatures ($\lesssim 100\,\mathrm{K}$), the mass is dominated by molecular gas, while at higher temperatures ($\gtrsim 20000,\mathrm{K}$), the ionized gas becomes the dominant component.
@@
Figure 8 shows the distributions of LOS velocity dispersions obtained for different temperature thresholds, for the simulations SN-1.5, TIGRESS-MHD4pc, TURB-7, and TURB-10. Varying the adopted temperature range produces only minor changes in the resulting dispersions. The relative order of the simulations remains unchanged, and the mean values shift by a few tenths of kilometers per second at most. In particular, SN-driven models remain systematically below the observational reference, even when molecular gas is excluded or a fraction of the ionized gas is included, while the large-scale forcing runs continue to exceed it. Our conclusions therefore do not depend sensitively on the adopted temperature range used to approximate the neutral atomic component.